# Nonlinear Model Order Reduction of Dynamical Systems in Process Engineering: Review and Comparison

### v1

@@ 4.1.2.  Model aggregation
In order to eliminate the potential IR in COMP ROMs and improve model sparsity, [168] presents the “aggregation method” (AGG). AGG is a heuristic modification of SPT for 1D distributed systems. Instead of transforming the FOM into the canonical form (12), the authors multiply (only) the left-hand side of the FOM by a matrix $\tilde{\mathbf{H}}=\mathrm{diag}(\tilde{h}_{1},\tilde{h}_{2},...,\tilde{h}_{n_{x}})$, $\tilde{h}_{i}\geq 0$. Thereby, the time constants of individual stage dynamics are artificially increased or reduced. $\tilde{\mathbf{H}}$ is a hyperparameter and commonly $\mathrm{trace}(\tilde{\mathbf{H}})\approx n_{x}$. Let $\mathbf{\Omega}$ be the $n_{x}\times n_{x}$ permutation matrix such that $\mathbf{\Omega}\tilde{\mathbf{H}}=\big{[}\begin{smallmatrix}\mathbf{H}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{smallmatrix}\big{]}$, where $\mathbf{H}=\mathrm{diag}(h_{1},h_{2},...,h_{n_{z}})$, $h_{i}>1$. The resulting ROM reads:
@@
A systematic comparison of AGG to COMP is given in [169]. By contrasting Eqs. 13 and 27, we notice that the manifolds on which the ROMs evolve are identical if $\bm{f}_{2}$ coincide. Hence, both methods may realize a different ROM living on the same manifold $\mathcal{M}$.
@@ 4.1.3.  Collocation method
Motivated by the collocation method for solving ODEs and PDEs [170, 171], the studies [8, 172] adapt the collocation method to MOR of spatially discrete (staged) problems. The underlying assumption is a polynomial coherence of the differential variables of the FOM. Then, a ROM is obtained by combining polynomial interpolation and a number of discrete cells (collocation points), e.g., finite volume elements or column stages. However, the collocation points do not necessarily correspond to cells in the FOM. Intuitively, the number and position of collocation points depend on the system, i.e., whether a low-order polynomial is able to capture the underlying pattern (slow manifold). The positions of the collocation points are typically chosen a-priori as the roots of Legendre, Hahn, or Jacobi polynomials and not further adapted online [173, 174]. In other words, the spatial collocation points are determined once and for all, i.e., do not change over time.
@@ 4.2.1.  Distillation columns
Detailed distillation column models typically include “stage-by-stage” differential energy, mass, and possibly momentum balances [175]. Common simplifying assumptions are ideally mixed stages with no diffusion effects, thermodynamic vapor-liquid equilibrium, and quasi-stationary pressure and enthalpy dynamics of both phases [86, 82]. These assumptions rely on time-scale assumptions and may be regarded as a heuristic application of SPT, rendering such mechanistic models already reduced. However, these models are often too expensive for online applications and therefore require MOR [176, 177].
@@
Using SPT, [85] derives a widely applicable COMP model form and shows a higher accuracy over COL. The COMP ROMs therein combine dynamic compartment balances and quasi-stationary stage models. The separation stage whose differential balances are replaced by compartment balances is termed “sensitivity stage”. The appearance of a non-physical IR depends on the number of compartments and the position of the sensitivity stages [87, 180]. However, avoiding an IR may limit the degree of order reduction. On the other hand, the particular structure of COMP models, featuring large sets of neighboring quasi-stationary stage models, enables hyperreduction via steady-state shortcut methods, e.g., FUG shortcut [46] or group methods [181, 182]. Typically, these shortcuts have fewer parameters than universal regressors (e.g., ANNs) and are less prone to overfitting. On the other hand, ANNs can reach almost arbitrary precision [183].
In steady operation, a distillation column exhibits a set of coherent composition patterns, which can be analytically derived under simplifying assumptions [184, 185]. These analytical solutions are wave-shaped profiles, and therefore called “nonlinear wave propagation model” (NWPM). [186] proposed a semi-empirical ROM for multi-species distillation columns, by combining NWPM equations, global column balances, and a boundary equilibrium assumption. The resulting NWPM-ROM is a compartment model (a non-physical IR is also observed here). In [187], this ROM was extended to account for effects due to non-constant holdup.
The AGG method was originally developed for distillation columns [168]. Like COMP models, AGG ROMs involve clusters of quasi-stationary trays that can be simplified using shortcut models. Different from COMP, the AGG ROM states may be associated with actual stages in a distillation column. While avoiding a non-phsical IR, the AGG states usually converge more slowly to the FOM states after a change of system inputs [188].
Besides the specialized MOR methods, general-purpose MOR methods have been applied to distillation column models, e.g., using POD-Galerkin [61, 189], standard SPT [190], KW [142, 120], and DMDc [191]. However, as opposed to these general-purpose methods, the specialized methods, COMP (including NWPM), AGG, and COL, build on the intrinsic system structure, providing generic low-order models that can be written (almost) independently of a FOM. In particular, these specialized methods can be pre-configured for generic columns with similar stage models and in many cases steady-state data is sufficient to configure these ROMs for a specific column. Moreover, COMP and AGG may access an extra set of hyperreduction methods in the form of steady-state shortcuts as described above.
@@ 4.2.2.  Heat exchangers
As an alternative to MOR, simplifying assumptions enable a heuristic degeneration or lumping of differential equations. Common assumptions are (i) a time scale separation between energy dynamics of fluids and walls and (ii) for parallel flow HX, a uniform “common wall” temperature in each cross section perpendicular to the principal flow axis [192]. Common examples of lumped parameter models are $\varepsilon$-NTU, P-NTU and LMTD [193]. However, these are typically only valid at stationary points and under simplified assumptions on geometry, heat transfer, and thermodynamics.
@@
Applications of general-purpose MOR methods to HX include POD-Galerkin [197, 198, 199], DMD [200], and POD combined with ANNs/RNNs [201]. Furthermore, both classical collocation and Galerkin projection are standard methods to approximate low-order solutions of HX PDE models [202, 193]. In addition, all three specialized approaches (COMP, AGG, COL) have been applied to spatially discretized dynamical models of 1D distributed HX. Specifically, [179, 203] use the collocation method to reduce finite volume models of HXs. Lumped COMP models of HX are presented in [204, 205, 206]. Also, the reduced modeling of HX with phase change (evaporators or condensers) is often accomplished using the “moving boundary model” (MBM) [207], which is a COMP model. The MBM uses compartments with flexible boundaries to account for spatially moving phase regions and can exhibit a non-physical IR [208]. Finally, AGG is applied to HX in [209].
Similar to separation columns, there exist steady-state shortcut models and analytic solutions to replace steady-state equations in COMP and AGG models. Common lumped parameter models are reviewed in [193]. Refs. [194] presents an analytic solution to the stationary multi-stream HX equations under simplifying assumptions. Further, HX networks may be used as a surrogate model of multi-stream HX [210, 211].
@@ 4.2.3.  Reactors
A large set of works has been devoted to constructing the invariant manifolds by approximately solving the invariance equations of chemical kinetic systems with non-obvious slow and fast partitioning. These methods employ system linearization [212], computational singular perturbation [217], thermodynamic dissipativity and symmetry [218] relation graphs [219, 220], or sensitivity analysis and numerical optimization [221, 222, 223]. Extensive reviews are given in [224, 225, 226]. Recently, ML methods have been increasingly applied to reduce kinetic systems [227, 228, 229, 230].
For spatially distributed reaction systems, e.g., reaction diffusion problems and tube reactors, the application of POD-Galerkin method [231, 232, 215, 233], nonlinear-Galerkin method [234, 13, 17], manifold-Galerkin method [36], and lift-and-reduce techniques [118] have been reported. As shown in [36], nonlinear subspace methods can significantly outperform linear basis methods in the very-low-order region. The classical collocation method is applied, e.g., in [235, 10, 236], to obtain low-order solutions of respective PDEs models. Moreover, an overview of compartment modeling of distributed chemical reactors is found in [237].
For integrated reaction-separation systems (with recycle), the works [238, 239] performed a time scale analysis by means of SPT to derive ROMs for control. Along the same lines, [196] investigated integrated reaction-heat exchange networks. Ref. [60] compared the POD-Galerkin and POD-Res method on an integrated reaction-separation system, finding similar ROM accuracy but higher CPU costs of POD-Res ROMs.
@@ 5.  Case study
We compare the performance of MOR methods on an industrial process by studying the ASU presented in [240]. Clearly, comparing all the methods reviewed in Sections 3 and 4 would exceed the scope of this article. Hence, we examine a subset of methods listed in Table 1. Specifically, we compare the MOR approaches (a-g) and (j). Additionally, we consider the two tailored methods COMP and AGG from Section 4.
@@
We briefly comment on the method selection. Despite successful applications in the literature, we neither investigate COL nor NWPM, because low-order COL models are known to exhibit a significant steady-state offset [85] and NWPM is a special case of COMP. We refer to [179] and [187] for applications of COL and NWPM to the ASU considered herein, respectively. Furthermore, we do not investigate OPI (h) due to the intricate “pen-and-paper” lifting [146]. Of the many possible approaches (i) and (j) reviewed in Section 3.2.4, we choose the MFL-ANN method [159] due to conceptual simplicity and low implementation efforts. Further, class (i) is theoretically included in (j) and therefore not investigated separately. In particular, we expect all single-stage methods in Section 3.2.4 to reach a comparable accuracy. Finally, recall that POD-Res and COMP represent data-driven and heuristic variants of SPT, respectively. In this regard, our case study covers SPT. As discussed in Section 1, we focus on MOR and do not investigate a potential hyperreduction and CPU times.
@@
Fig. 4 illustrates the ASU under investigation, which produces gaseous nitrogen and is built of a main air compressor (MAC) with precooling, three multi-stream HXs, two turbines (T1, T2), a high-pressure rectifying column (HPC) with 30 trays, and an integrated reboiler condenser unit (IRC). The ASU has four manipulated variables (MVs), which are the molar flow rate of air $F_{\mathrm{mac}}$, the molar flow fraction to the turbine $\zeta_{\mathrm{turb}}$, the reflux fraction at the column top $\zeta_{\mathrm{cond}}$, and the purge stream $F_{\mathrm{drain}}$.
@@
The ASU is open-loop unstable due to the integrator behavior of the liquid molar holdup $M_{\mathrm{irc}}$ of the IRC. For unstable processes, data sampling is more challenging, especially if only a small subregion of the state space is relevant for ROM application. However, most chemical plants feature a stabilizing base layer as part of the automation hierarchy, so stability is not an issue [241]. For simplicity, we thus assume that the IRC is equipped with an ideal inventory controller that manipulates $F_{\mathrm{drain}}$ to stabilize $M_{\mathrm{irc}}$ at its nominal value. Additionally, we fix $\zeta_{\mathrm{turb}}$ to its nominal value. Consequently, we have two variable inputs $\bm{u}=[F_{\mathrm{mac}},\zeta_{\mathrm{cond}}]^{\mathrm{T}}$.
@@
For our purpose, the selected ASU features some characteristic properties of industrial process systems, such as variable coupling and nonlinear process response. Moreover, we regard the ASU as a good compromise between the complexity of the process system and the complexity of the MOR problem. However, the selected process does not incorporate a chemical reactor and has few inputs. The investigation of MOR for complex processes including reaction and many MVs is indispensable and should be undertaken in future studies.
Fig. 4: Process flowsheet of the ASU: Heat exchangers (HX), high-pressure column (HPC), reboiler tank (IRC), compressor (MAC), and turbines (T). Blue circles indicate inputs $\bm{u}(t)$.
@@ 5.1.  Full-order model
We consider the three main species of air nitrogen, argon, and oxygen. In contrast, we do not model carbon dioxide, hydrocarbons, or other residual chemical species here. The FOM of the ASU comprises material and energy balances of all process units as well as constitutive equilibrium-based thermodynamic equations. The HPC is modeled in a tray-to-tray fashion assuming thermal equilibrium with constant relative volatilities, constant specific heat capacities of all species, and ideal mixture (no excess properties). Both the HPC bottom tray and the IRC reboiler are modeled as special column trays with an extra feed stream and fixed liquid outflow, respectively. The IRC condenser is a total condenser with negligible material holdup and modeled using lumped quasi-stationary balance equations. We apply a logarithmic transformation of molar fractions to support the extraction of composition patterns, i.e., the logarithmic molar fractions are the respective differential states. The HX models feature energy balances for the metal wall and quasi-stationary balances for the fluids, assuming a single stream layer each, 1D spatial discretization with $n_{\mathrm{HX}}$ cells per stream, and the common wall assumption. Following [240], we set $n_{\mathrm{HX1}}=25$, $n_{\mathrm{HX2}}=25$, and $n_{\mathrm{HX3}}=1$. Lastly, first-order low-pass filters account for the lagged vapor flow rate in HX and HPC.
In total, the FOM comprises $n_{u}=2$ inputs and $n_{x}=176$ differential states, and is implemented in Modelica. The admissible input set $\mathcal{U}$ is a box defined by $F_{\mathrm{mac}}\in[30\,\mathrm{mol/s},50\,\mathrm{mol/s}]$ and $\zeta_{\mathrm{cond}}\in[0.51,0.54]$. As the initial state $\bm{x}_{0}$, we use the nominal steady state corresponding to $F_{\mathrm{mac}}=52.2\,\mathrm{mol/s}$, $\zeta_{\mathrm{cond}}=0.535$, $\zeta_{\mathrm{turb}}=0.805$, and $M_{\mathrm{irc}}=25\,\mathrm{kmol}$. At $\bm{x}_{0}$, the ASU produces 25 mol/s nitrogen of 100 ppm impurity grade.

### v2

@@ 4.1.2. Model aggregation
In order to eliminate the potential inverse response in COMP ROMs and improve model sparsity, [163] presents the aggregation method (AGG). AGG is a heuristic modification of SPT for 1D distributed systems. Instead of transforming the FOM into the canonical form (13), the authors multiply (only) the left-hand side of the FOM by a matrix $\tilde{\mathbf{H}}=\mathrm{diag}(\tilde{h}_{1},\tilde{h}_{2},...,\tilde{h}_{n_{x}})$, $\tilde{h}_{i}\geq 0$. Thereby, the time constants of individual stage dynamics are artificially increased or reduced. $\tilde{\mathbf{H}}$ is a hyperparameter and commonly selected such that $\mathrm{trace}(\tilde{\mathbf{H}})\approx n_{x}$. Let $\mathbf{\Omega}$ be the $n_{x}\times n_{x}$ permutation matrix such that $\mathbf{\Omega}\tilde{\mathbf{H}}=\big[\begin{smallmatrix}\mathbf{H}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{smallmatrix}\big]$, where $\mathbf{H}=\mathrm{diag}(h_{1},h_{2},...,h_{n_{z}})$, $h_{i}>1$. The resulting ROM reads:
@@
A systematic comparison of AGG to COMP is given in [168]. By contrasting (14) and (28), we notice that the manifolds on which the ROMs evolve are identical if $\bm{f}_{2}$ coincide. Hence, both methods may realize a different ROM living on the same manifold $\mathcal{M}$.
@@ 4.1.3. Collocation method
Motivated by the collocation method for solving ordinary and partial differential equations [102, 71], the studies [299, 49] adapt the collocation method to MOR of spatially discrete problems. The underlying assumption is a polynomial coherence of the differential variables of the FOM. Then, a ROM is obtained by combining polynomial interpolation and a number of discrete cells (collocation points), e.g., finite volume elements or column stages. However, the collocation points do not necessarily correspond to cells in the FOM. Intuitively, the number and position of collocation points depend on the system, i.e., whether a low-order polynomial is able to capture the underlying patterns characterizing the slow manifold. The positions of the collocation points are typically chosen a-priori as the roots of Legendre, Hahn, or Jacobi polynomials and not further adapted online [274, 277]. In other words, the spatial collocation points are determined once and for all, i.e., do not change over time.
@@ 4.2.1. Distillation columns
Detailed distillation column models typically include stage-by-stage differential energy, mass, and possibly momentum balances [126]. Common simplifying assumptions are ideally mixed stages with no diffusion effects, thermodynamic vapor-liquid equilibrium, and quasi-stationary pressure and enthalpy dynamics of both phases [234, 244]. These assumptions rely on time-scale assumptions and may be regarded as a heuristic application of SPT, rendering such mechanistic models already reduced. However, these models are often too expensive for online applications and therefore require MOR [1, 259].
@@
Using SPT, [20] derives a widely applicable COMP model form and shows a higher accuracy over COL. The COMP ROMs therein combine dynamic compartment balances and quasi-stationary stage models. The separation stage whose differential balances are replaced by compartment balances is termed sensitivity stage. The appearance of a non-physical inverse response depends on the number of compartments and the position of the sensitivity stages [125, 127]. However, avoiding an inverse response may limit the degree of order reduction. On the other hand, the particular structure of COMP models, featuring large sets of neighboring quasi-stationary stage models, enables hyperreduction via steady-state shortcut methods, e.g., FUG shortcut [58] or group methods [135, 60]. Typically, these shortcuts have fewer parameters than universal regressors (e.g., ANNs) and are less prone to overfitting. On the other hand, ANNs can reach almost arbitrary precision [254].
In steady operation, a distillation column exhibits a set of coherent composition patterns, which can be analytically derived under simplifying assumptions [185, 128]. These analytical solutions are wave-shaped profiles, and therefore called nonlinear wave propagation model. Ref. [141] proposed a semi-empirical ROM for multi-species distillation columns, by combining wave propagation equations, global column balances, and a boundary equilibrium assumption. The resulting ROM is a compartment model (a non-physical inverse response is also observed here). In [43], this ROM was extended to account for effects due to non-constant holdup.
The AGG method was originally developed for distillation columns [163]. Like COMP models, AGG ROMs involve clusters of quasi-stationary trays that can be simplified using shortcut models. Different from COMP, the AGG ROM states may be associated with actual stages in a distillation column. While avoiding a non-physical inverse response, the AGG predictions usually converge more slowly to the FOM states after a change of system inputs [167].
Besides the specialized MOR methods, general-purpose MOR methods have been applied to distillation column models, e.g., using POD-Galerkin [99, 113], standard SPT [154], KW [262, 260], and DMDc [231]. However, as opposed to these general-purpose methods, the specialized methods, COMP, nonlinear wave propagation, AGG, and COL, build on the intrinsic system structure, providing generic low-order models that can be written (almost) independently of a FOM. In particular, these specialized ROMs can be pre-configured for generic columns with similar stage models and in many cases steady-state data is sufficient to customize these ROMs for a specific column. Moreover, COMP and AGG may access an extra set of hyperreduction methods in the form of steady-state shortcuts as described above.
@@ 4.2.2. Heat exchangers
As an alternative to MOR, simplifying assumptions enable a heuristic degeneration or lumping of differential equations. Common assumptions are (i) a time scale separation between energy dynamics of fluids and walls and (ii) for parallel flow HX, a uniform common wall temperature in each cross section perpendicular to the principal flow axis [228]. Common examples of lumped parameter models are $\varepsilon$-NTU, P-NTU and LMTD [264]. However, these are typically only valid at stationary points and under simplified assumptions on geometry, heat transfer, and thermodynamics.
@@
Applications of general-purpose MOR methods to HX include POD-Galerkin [302, 50, 178], DMD [303], and POD combined with ANN predictors [166]. Furthermore, both classical collocation and Galerkin projection are standard methods to approximate low-order solutions of partial differential equations [106, 264]. In addition, all three specialized approaches (COMP, AGG, COL) have been applied to spatially discretized dynamical models of 1D distributed HX. Specifically, [41, 233] use the collocation method to reduce finite volume models of HXs. Lumped COMP models of HX are presented in [188, 306, 192]. Also, the reduced modeling of HX with phase change (evaporators or condensers) is often accomplished using the moving boundary model [132], which is a COMP model. The moving boundary model uses compartments with flexible boundaries to account for spatially moving phase regions and can exhibit a non-physical inverse response [285]. Finally, AGG is applied to HX in [169].
Similar to separation columns, there exist steady-state shortcut models and analytic solutions to replace steady-state equations in COMP and AGG models. Common lumped parameter models are reviewed in [264]. The work [173] presents an analytic solution to the stationary multi-stream HX equations under simplifying assumptions. Further, HX networks may be used as a surrogate model of multi-stream HX [305, 111].
@@ 4.2.3. Reactors
A large set of works has been devoted to constructing invariant manifolds by solving the invariance equations of chemical kinetic systems with non-obvious slow and fast partitioning. These methods employ system linearization [180], computational singular perturbation [157], thermodynamic dissipativity and symmetry [88] relation graphs [172, 222], or sensitivity analysis and numerical optimization [61, 225, 28]. Extensive reviews are given in [87, 48, 272]. Recently, machine learning has been increasingly applied to reduce kinetic systems [133, 294, 137, 216]. Interestingly, the chemical reaction network formalism has also been applied as mathematical framework to facilitate quadratization of generic ODEs [115, 65].
For spatially distributed reaction systems, e.g., reaction diffusion problems and tube reactors, the application of POD-Galerkin method [214, 152, 123, 7], nonlinear-Galerkin method [69, 92, 269], manifold-Galerkin method [159], and lift-and-reduce techniques [149, 21] have been reported. As shown in [159], nonlinear subspace methods can significantly outperform linear basis methods in the very-low-order region. The classical collocation method is applied, e.g., in [105, 80, 161], to obtain low-order solutions of respective partial differential equations. Moreover, an overview of compartment modeling of distributed chemical reactors is found in [97].
For integrated reaction-separation systems (with recycle), the works [51, 153] performed a time scale analysis by means of SPT to derive ROMs for control. Along the same lines, [13] investigated integrated reaction-HX networks. Ref. [256] compared the POD-Galerkin and POD-Res method on an integrated reaction-separation system, finding similar ROM accuracy but higher CPU costs of POD-Res ROMs.
@@ 5. Case study
We compare the performance of MOR methods on an industrial process by studying the air separation unit presented in [44]. Clearly, comparing all the methods reviewed in Sections 3 and 4 would exceed the scope of this article. Hence, we examine a subset of methods listed in Table 2. Specifically, we compare the MOR approaches (a-g) and (j). Additionally, we consider the two tailored methods COMP and AGG from Section 4.
@@
We briefly comment on this choice of methods. Despite successful applications in the literature, we investigate neither COL nor nonlinear wave propagation models, because the former are known to exhibit a significant steady-state offset [20] and the latter are a special case of COMP. We refer to [41, 43] for applications of these methods to the process considered herein. As discussed in Section 3.2.3, we regard KW and DMDc as a variant of operator inference (h) and do not investigate other forms herein. Of the many possible approaches (i) and (j) reviewed in Section 3.2.4, we choose the MFL-ANN method [187] due to conceptual simplicity and low implementation efforts. Further, class (i) is theoretically included in (j) and therefore not investigated separately. In particular, we expect all single-stage methods in Section 3.2.4 to reach a comparable accuracy. Finally, recall that POD-Res and COMP represent data-driven and heuristic variants of SPT, respectively. In this regard, our case study covers SPT. As discussed in Section 1, we focus on MOR and do not investigate a potential hyperreduction and CPU times.
@@
Figure 3 illustrates the process under investigation, which produces gaseous nitrogen and is built of a main air compressor with precooling, three multi-stream HXs, two turbines, a high-pressure rectifying column with 30 trays, and an integrated reboiler-condenser unit. The unit has four manipulated input variables, which are the molar flow rate of air $F_{\mathrm{mac}}$, the molar flow fraction to the turbine $\zeta_{\mathrm{turb}}$, the reflux fraction at the column top $\zeta_{\mathrm{cond}}$, and the purge stream $F_{\mathrm{drain}}$.
@@
The system is open-loop unstable due to the integrator behavior of the liquid molar holdup $M_{\mathrm{irc}}$ of the reboiler tank. For unstable processes, data sampling is more challenging, especially if only a small subregion of the state space is relevant for ROM application. However, most chemical plants feature a stabilizing base layer as part of the automation hierarchy, so stability is not an issue [63]. For simplicity, we thus assume that the reboiler tank is equipped with an ideal inventory controller that manipulates $F_{\mathrm{drain}}$ to stabilize $M_{\mathrm{irc}}$ at its nominal value. Additionally, we fix $\zeta_{\mathrm{turb}}$ to its nominal value. Consequently, we have two variable inputs $\bm{u}=[F_{\mathrm{mac}},\zeta_{\mathrm{cond}}]^{\mathrm{T}}$.
@@
For our purpose, the selected process features some characteristic properties of industrial process systems, such as variable coupling and nonlinear process response. Moreover, we regard the process as a good compromise between the complexity of the process system and the complexity of the MOR problem. However, the selected process does not incorporate a chemical reactor and has few inputs. The investigation of MOR for complex processes including reaction and many input variables is indispensable and should be undertaken in future studies.
Fig. 3: Schematic process flowsheet of the investigated air separation unit. Blue circles indicate the system inputs $\bm{u}(t)$.
@@ 5.1. Full-order model
We consider the three main species of air nitrogen, argon, and oxygen. In contrast, we do not model carbon dioxide, hydrocarbons, or other residual chemical species here. The FOM comprises material and energy balances of all process units as well as constitutive equilibrium-based thermodynamic equations. The column is modeled in a tray-to-tray fashion assuming thermal equilibrium with constant relative volatilities, constant specific heat capacities of all species, and ideal mixture (no excess properties). Both the column bottom tray and the reboiler are modeled as special column trays with an extra feed stream and fixed liquid outflow, respectively. The condenser is a total condenser with negligible material holdup and modeled using lumped quasi-stationary balance equations. We apply a logarithmic transformation of molar fractions to support the extraction of composition patterns, i.e., the logarithmic molar fractions are the respective differential states. The HX models feature energy balances for the metal wall and quasi-stationary balances for the fluids, assuming a single flow chanel for each stream, 1D spatial discretization with $n_{\mathrm{HX}}$ cells per stream, and the common wall assumption. Following [44], we set $n_{\mathrm{HX1}}=25$, $n_{\mathrm{HX2}}=25$, and $n_{\mathrm{HX3}}=1$. Lastly, first-order low-pass filters account for the lagged vapor flow rate in heat exchangers and column.
In total, the FOM comprises $n_{u}=2$ inputs and $n_{x}=176$ differential states, and is implemented in Modelica. Although $n_{x}$ may appear low compared to some structural mechanic or fluid dynamics problems, the system investigated here constitutes a representative example of model reduction in the process control field, cf. [259, 233, 62]. The admissible input set $\mathcal{U}$ is a box defined by $F_{\mathrm{mac}}\in[30\,\mathrm{mol/s},50\,\mathrm{mol/s}]$ and $\zeta_{\mathrm{cond}}\in[0.51,0.54]$. As the initial state $\bm{x}_{0}$, we use the nominal steady state corresponding to $F_{\mathrm{mac}}=52.2\,\mathrm{mol/s}$, $\zeta_{\mathrm{cond}}=0.535$, $\zeta_{\mathrm{turb}}=0.805$, and $M_{\mathrm{irc}}=25\,\mathrm{kmol}$. At $\bm{x}_{0}$, the process produces 25 mol/s nitrogen of 100 ppm impurity grade.